# Monitoring of water volume in a porous reservoir using seismic data: Validation of a numerical model with a field experiment

### v2

@@ 1 Introduction
Groundwater aquifers are facing unprecedented threats, with levels decreasing at alarming rates, often more than one meter per year in some areas—leading to significant long-term reductions. As a result, surface water flows that were previously sustained by groundwater are becoming seasonal or disappearing completely [9]. To ensure sustainable water extraction, it is crucial to have a better understanding of the location and extent of groundwater resources. In response to these challenges, this study focuses on validating a numerical model using seismic data from a small-scale experimental setup with emphasis to estimate water volume in a porous reservoir at various water table levels, i.e. water volumes. Neural networks have been increasingly applied to enhance the analysis of seismic data, for a recent review, see [18]. By using neural network techniques, we aim to directly estimate water volume from seismic responses, avoiding the traditional need for separate determinations, for example, of reservoir water table level and porosity.
Geophysical methods, including seismic techniques, are commonly used tools in the early stages of groundwater exploration and for ensuring sustainable extraction strategies [21, 10, 8]. They offer a cost-effective alternative to drilling, providing laterally continuous data across vast areas, by utilizing variations in material properties to detect subsurface features. Seismic methods, in particular, are well-suited for locating and monitoring water resources due to the higher seismic velocities exhibited by saturated materials compared to unsaturated ones [12]. The high resolution of seismic methods, both horizontally and vertically, enables the creation of detailed subsurface feature mapping.
@@ 2.2 Seismic measurements
In our experiments, a metallic rod weight was dropped on a steel plate and served as a source of seismic waves, with an electric brake mechanism to control the rod’s release and prevent multiple hits. The experiments involved shots at 13 specific locations and were conducted at three different drop heights ($\rm{dh}$), consisting of 5, 10, and 15 cm elevations. Three different heights were selected to test the signal amplitude clipping threshold in the measurement device.
@@ 3.1 Synthetic model of the Laukaa test site
The applied synthetic model shown in Fig. 2 replicates the geometry of the Laukaa test site. It is a box with a length of 31.5 m, a width of 16.2 m, and a height of 2.75 m. In addition, the corners of the geometry are rounded as shown in the graph. The maximum length and width of the air-saturated zone are 29.5 m and 14.2 m, respectively. Finally, the bottom profile of the water-saturated zone is a rectangle with a length of 23.5 m and a width of 8.2 m. The bottom is located at a depth of 2 m from the top surface. A water table divides the porous material into air-saturated and water-saturated subdomains. In the synthetic model, the water table level is allowed to move freely from $z=-25$ cm to $z=-1.2$ m.
@@
We assume a free boundary condition on the top surface while other boundaries are modeled as absorbing boundaries. In addition, we have a total of 57 receivers, which will capture the $y$- and $z$-directional solid velocity components $v_{\rm s}$ and $w_{\rm s}$. Exact receiver locations were accurately surveyed using a commercially available differential (DGPS) system, and these locations are used in the numerical modeling. One extra receiver, illustrated by a green color, is located next to line 1 and is used as a reference point. A more detailed discussion of the reference point is given later in Section 4.1.
Although the measurement data acquisition involved 13 shot locations, the water volume estimation is based on data from a single source location only, as used in the simulation study [13]. The location for the source was arbitrarily selected to be the closest to receiver 4 on Line 1. The seismic source is modeled as a vertical force pointing to the negative $z$-axis. Since the exact location of the source slightly varies between different measurements (distance from the closest geophone was measured with a ruler), the location was assumed to be uncertain in the numerical model when training the neural network. In practice, the source center location in the $(x,y)$-plane is randomized from $x\sim\mathcal{U}(4.4890,4.5890)$ m and $y\sim\mathcal{U}(-5.2142,-5.1142)$ m. Receivers and the source are placed on the ground surface.
@@ 3.2 Physical parameters
The volume of stored water in the reservoir can be calculated by multiplying the volume of the water-saturated domain by the porosity. In this paper, the amount of water is calculated only from the water-saturated zone that is located exactly under the array of receivers. In general, the accurate estimation of water volume is possible in water-saturated subdomain with the current problem setup, however, this cannot be principally assumed in more realistic aquifer models.
@@ 4.1 Noise model and source function normalization
where $\bm{m}$ contains all the physical and geometrical parameters of the model and $e$ accommodates additive noise components. The forward operator $\mathcal{A}$ is used to map the model parameters $\bm{m}$ to the seismic data vector, simulated using the coupled viscoelastic-poroviscoelastic material model by the DG method in three spatial dimensions [5, 13].
To approximate the measurement noise in the field measurements, we apply the following noise model
$X_{\ell}^{\textrm{noised}}=X_{\ell}+A|X_{\ell}|_{\max}\epsilon^{A}+B|X_{\ell}|\epsilon^{B},$ (5)
where $\epsilon^{A}$ and $\epsilon^{B}$ represent independent zero-mean Gaussian random variables. $|X_{\ell}|_{\max}$ is the maximum absolute value of the $\ell$’th sample. The two noise components correspond to additive white noise and amplitude-related noise, respectively, contributing to a diverse range of noise levels. By adjusting $A$ and $B$ within the intervals $[0.05,2]$ per cent and $[0,2]$ per cent consecutively, wide noise variations are introduced.
@@
where $F_{\{\Upsilon,\Upsilon_{\rm ref},X,X_{{\rm ref}}\}}$ represents the data in the frequency domain obtained through the Fourier transform. The subscript ${\rm ref}$ denotes the data corresponding to the reference point. $\hat{e}$ is the noise in the frequency domain formulation after normalization, $\omega_{\ell}$ is the frequency at the $\ell$th index, and $N_{f}$ indicates the total number of frequencies. However, (7) can become ill-conditioned when the denominator approaches zero. To address this issue, we adopt the Wiener filtering approach for regularization, see, for example, [22].
@@ 4.2 Examples of real and synthetic seismic data
Figure 3 presents a sample shot gather from the real dataset, displaying the $v_{\rm s}$ and $w_{\rm s}$ velocity components along with the reference records as an example in the top right corner. The bottom row shows the real and imaginary parts of the Fourier transformed and source function normalized records, which serve as inputs to the neural network. For the shown data, water table level was at $-88.7$ cm.
Figure 3: An example of the measured data. The figures on the left show the $v_{\rm s}$ component, and the figures on the right the $w_{\rm s}$ component. The velocity components on the reference point are shown to illustrate the data behaviour.
Figure 4 shows an example of the simulated data. Here the water table level was at $-72.10$ cm and the noise level parameters were set to $A=1$ per cent and $B=1$ per cent. The discrepancy between measured and synthetic seismograms in these figures is due to the random selection of material parameters employed in generating the synthetic data.
Figure 4: An example of the synthetic data. The pictures on the left show the $v_{\rm s}$ component, and the figures on the right the $w_{\rm s}$ component. The reference data traces on the top panel are shown for the noisy and noiseless data for both velocity components.
@@ 4.3 Neural networks
We generated 15,000 training and 3,000 validation samples using fifth-order polynomials and the source wavelet (1) with the DG solver. The inviscid material model-based element size criterion is approximately 1.9 elements per shortest wavelength for the samples in the validation dataset and 2.0 for training dataset samples.
@@
A fully connected neural network was employed in this study to determine the water volume using seismic data. We down-sample the synthetic data to the sampling frequency of 4 kHz, then five copies of clean training and validation databases are corrupted with Gaussian noise according to the noise model in (5). The time domain data are transformed to the frequency domain and the corresponding real and imaginary components are used as inputs to the network. We select 21 frequencies between 30 to 90 Hz, see Figs. 3 and 4 as examples.
@@
The training uses the TensorFlow [2] and Keras [4] libraries. We use the Adam optimizer [14] and the RandomSearch algorithm of the Keras Tuner library [19] which maximizes the model’s behavior given the training and validation datasets by testing different hyperparameters in different random combinations. We allow Keras Tuner to randomize the activation function, learning rate, number of hidden layers, number of neurons per hidden layer, and L2 regularization penalty factor. Our choices for activation function, are “relu”, “sigmoid”, “tanh”, “selu”, “swish”, or “LeakyReLU” and number of hidden layers from one to six. The number of neurons varies from 100 to 5,000, and for the learning rate, we allow Keras Tuner to choose from [1e-3, 1e-4, 1e-5], and L2 regularization penalty from [1e-5, 1e-6, 1e-7, 1e-8]. For all networks studied in this work, the batch size is set to 256.
@@
In this study, the best-performing network is selected according to mean absolute error (MAE) with validation. After testing the different hyperparameter combinations explained above, the network is made up of five hidden layers with 2,570 (layer 1), 3,920 (layer 2), 3,360 (layer 3), 2,730 (layer 4), and 3,400 (layer 5) neurons, a LeakyRelu activation function, a learning rate of 1e-5, and L2 regularization penalty factor of 1e-5. Linear activation is used for the output layer, and early stopping is also activated both in the hyperparameter tuning and actual training phases.
@@ 5.1 Predictions of water volume
  • Field Measurements: Field data have been assessed at seven distinct water table levels, ranging from a depth of -31.3 cm to -88.7 cm. For each water level, we collected three repeated measurements from three different drop heights. These lead to a dataset consisting of a total of $3\times 3\times 7=63$ samples. Using our knowledge of the actual water table level and the geometry of the sand pool, we can calculate the volume of the water-saturated zone. Additionally, based on a previous study [20], we assume the nominal porosity of the material to be 35 per cent.
@@
The Keras Tuner-optimized network architecture undergoes ten training runs, and the final water volume per sample result is determined as the average of these ten estimates. Figure 5 presents a comparison between the estimated (average) water volumes and their corresponding true values. The results for all drop heights are shown with different colors, red color denotes the repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). To get a crude approximation for the uncertainty, we assumed that the porosity value in the field measurements database is uncertain in a sense that we assumed $\phi_{\rm{true}}\in[0.95,1.05]\cdot 35$ per cent, that can be used to compute the error bars shown for each estimate with real data. The figure shows also the estimates for the synthetic dataset. These results demonstrate the potential of using proposed neural network based approach to recover the water volume.
Figure 5: Left: Estimated volumes of water as a function of true value. Right: Histogram of relative prediction errors.
@@ 5.2 SHAP analysis
Next, we constructed two new receiver configurations and trained the neural network model for the field measurements based on SHAP values. For the first configuration, we selected ten receivers having the largest SHAP values, and for the second configuration, we randomly selected ten receivers from the full sensor array (see bottom panel of Fig. 6). The KerasTuner optimized network with SHAP analysis-based receiver selection consists three hidden layers with 3,490 (layer 1), 3,780 (layer 2), and 2,580 (layer 3) neurons, a LeakyRelu activation function, a learning rate of 1e-4, and L2 regularization penalty factor of 1e-6. Similarly for the randomly selected receiver selection lead to optimized network with two hidden layers with 3,490 (layer 1) and 2,580 (layer 2) neurons, a LeakyRelu activation function, a learning rate of 1e-4, and L2 regularization penalty factor of 1e-6.
@@
The NMB, MAE and RMSE are used as evaluation metrics to quantitatively analyze the estimation results for all three receiver configurations, see Table 4. The table shows that the full receiver array results in the most accurate measures.
@@
Receiver array setup Database NMB (%) MAE (m3) RMSE

### v3

@@ 1 Introduction
Groundwater aquifers are facing unprecedented threats, with levels decreasing at alarming rates, often more than one meter per year in some areas—leading to significant long-term reductions. As a result, surface water flows that were previously sustained by groundwater are becoming seasonal or disappearing completely [9]. To ensure sustainable water extraction, it is crucial to have a better understanding of the location and extent of groundwater resources. In response to these challenges, this study focuses on validating a numerical model using seismic data from a small-scale experimental setup with emphasis to estimate water volume in a porous reservoir at various water table levels. Neural networks have been increasingly applied to enhance the analysis of seismic data, for a recent review, see [19]. By using neural network techniques, we aim to directly estimate water volume from seismic responses, avoiding the traditional need for separate determinations, for example, of reservoir water table level and porosity.
Geophysical methods, including seismic techniques, are commonly used tools in the early stages of groundwater exploration and for ensuring sustainable extraction strategies [22, 11, 8]. They offer a cost-effective alternative to drilling, providing laterally continuous data across vast areas, by utilizing variations in material properties to detect subsurface features. Seismic methods, in particular, are well-suited for locating and monitoring water resources due to the higher seismic velocities exhibited by saturated materials compared to unsaturated ones [13]. The high resolution of seismic methods, both horizontally and vertically, enables detailed subsurface feature mapping.
@@ 2.2 Seismic measurements
In our experiments, a metallic rod weight was dropped on a steel plate and served as a source of seismic waves, with an electric brake mechanism to control the rod’s release and prevent multiple hits. The experiments involved shots at 13 specific locations and were conducted at three different drop heights ($\rm{dh}$), consisting of 5, 10, and 15 cm elevations. Three different heights were selected to test the signal amplitude clipping threshold at the nearest seismic receivers.
@@ 3.1 Synthetic model of the Laukaa test site
The applied synthetic model shown in Fig. 2 replicates the geometry of the Laukaa test site. It is a box with a length of 31.5 m, a width of 16.2 m, and a height of 2.75 m. In addition, the corners of the geometry are rounded as shown in the graph. The maximum length and width of the air-saturated zone are 29.5 m and 14.2 m, respectively. Finally, the bottom profile of the water-saturated zone is a rectangle with a length of 23.5 m and a width of 8.2 m. The bottom is located at a depth of 2 m from the top surface. A water table divides the porous material into air-saturated and water-saturated subdomains. When creating the data with the synthetic model, the water table level $z$ is randomized from $z\sim\mathcal{U}(-120,-25)$ cm.
@@
Our model includes a total of 57 receivers that record the solid velocity components in the horizontal ($y$) and vertical ($z$) directions, denoted by $v_{\rm s}$ and $w_{\rm s}$, respectively. These components represent the solid velocity in the model’s coordinate system (see Fig. 2). Notably, they do not correspond to isolated wave modes (e.g., P-, SV-, or SH-waves), but rather capture the full wavefield projected onto the model axes. As a result, the signals may contain contributions from multiple wave types, due to mode conversions and the complex nature of wave propagation in porous media.
Exact receiver locations were accurately surveyed using a commercially available differential GPS (DGPS) system and are used in the numerical modeling. One additional receiver, marked in green, is located near line 1 and serves as a reference point. A more detailed discussion of this reference point is provided in Section 4.1.
Although the measurement data acquisition involved 13 shot locations, the water volume estimation is based on data from a single source location only, as used in the simulation study [14]. The location for the source was arbitrarily selected to be the closest to receiver 4 on Line 1. The seismic source is modeled as a vertical force pointing to the negative $z$-axis. Since the exact location of the source slightly varies between different measurements (distance from the closest geophone was measured with a ruler), the location was assumed to be uncertain in the numerical model when training the neural network. In practice, to account for uncertainty in repositioning the source between measurements at different water table levels, we assume a 10 cm variability in the source location in both the $x$ and $y$ directions. This means that the source center location in the $(x,y)$-plane is randomized as $x\sim\mathcal{U}(4.4890,4.5890)$ m and $y\sim\mathcal{U}(-5.2142,-5.1142)$ m. Receivers and the source are placed on the ground surface.
@@ 3.2 Physical parameters
The volume of stored water in the reservoir can be calculated by multiplying the volume of the water-saturated domain by the porosity. In this paper, the amount of water is calculated only from the water-saturated zone that is located exactly under the array of receivers. The current problem setup, with its simplified geometric structure, allows for the potential estimation of water volume across the entire reservoir. However, this cannot be assumed to be directly applicable to more complex and realistic aquifer models.
For the Laukaa test case studied in this paper, the parameter distributions shown in Tables 1 and 2 are chosen to be representative of the geological properties observed in the region. However, it is important to note that while these ranges are designed to capture key variations, the neural network model may still perform well even if not all parameters are completely accounted for in the training data. This is because the network is able to learn relevant features from the available data, potentially allowing it to generalize and provide accurate estimates even in regions not explicitly covered by the training set. In this sense, the model can adapt to variations in parameters and still maintain robustness in its predictions [10].
@@ 4.1 Noise model and source function normalization
where $\bm{m}$ contains all the physical and geometrical parameters of the model and $e$ accommodates additive noise components. The forward operator $\mathcal{A}$ is used to map the model parameters $\bm{m}$ to the synthetic seismic data vector $X$, simulated using the coupled viscoelastic-poroviscoelastic material model by the DG method in three spatial dimensions [5, 14].
To simulate the noise of the field measurements, we apply the following noise model
$X_{\textrm{noised}}=X+A|X|_{\max}\epsilon^{A}+B|X|\epsilon^{B},$ (5)
where $\epsilon^{A}$ and $\epsilon^{B}$ represent independent zero-mean Gaussian random variables. $|X|_{\max}$ is the maximum absolute value of $X$. The two noise components correspond to additive white noise and amplitude-related noise, respectively, contributing to a diverse range of noise levels. By adjusting $A$ and $B$ within the intervals $[0.05,2]$ per cent and $[0,2]$ per cent, respectively, wide noise variations are introduced.
@@
where $F_{\{\Upsilon,\Upsilon_{\rm ref},X,X_{{\rm ref}}\}}$ represents the data in the frequency domain obtained through the Fourier transform. The subscript ${\rm ref}$ denotes the data corresponding to the reference point. $\hat{e}$ is the noise in the frequency domain formulation after normalization, $\omega_{\ell}$ is the frequency at the $\ell$th index, and $N_{f}$ indicates the total number of frequencies. Note that the model (7) is applied separately to both velocity components, $v_{\rm s}$ and $w_{\rm s}$, at each receiver. Since (7) may become ill-conditioned when the denominator approaches zero, we adopt a Wiener filtering approach for regularization, following the method applied in [23].
@@ 4.2 Examples of real and synthetic seismic data
Figure 3 presents a sample shot gather from the real dataset, displaying the horizontal $v_{\rm s}$ and vertical $w_{\rm s}$ velocity components along with the reference records as an example in the top right corner. The bottom row shows the real and imaginary parts of the Fourier transformed and source function normalized records, which serve as inputs to the neural network. For the shown data, water table level was at $-88.7$ cm.
Figure 3: An example of the measured data. The figures on the left show the horizontal $v_{\rm s}$ component, and the figures on the right the vertical $w_{\rm s}$ component. The velocity components on the reference point are shown to illustrate the data behaviour.
Figure 4 shows an example of the simulated data. Here the water table level was at $-72.10$ cm and the noise level parameters were set to $A=1$ per cent and $B=1$ per cent. The discrepancies between the measured and synthetic seismograms in these figures arise from the random selection of material parameters used to generate the synthetic data, as well as differences in the water table levels between the real and synthetic datasets.
Figure 4: An example of the synthetic data. The pictures on the left show the horizontal $v_{\rm s}$ component, and the figures on the right the vertical $w_{\rm s}$ component. The reference data traces on the top panel are shown for the noisy and noiseless data for both velocity components.
@@ 4.3 Neural networks
We generated 15,000 training and 3,000 validation samples using fifth-order polynomials and the source wavelet (1) with the DG solver. The values for the model input $\bm{m}$ were independently drawn from the propability distributions described in Sections 3.1 and 3.2 and used as inputs to the forward model $\mathcal{A}$ (see Eq. (7)). The inviscid material model-based element size criterion is approximately 1.9 elements per shortest wavelength for the samples in the validation dataset and 2.0 for training dataset samples.
@@
A fully connected neural network was employed in this study to determine the water volume using seismic data. By training on data with randomized model parameters $\bm{m}$ (see Eq. (7)), the network learns to focus on the relevant seismic signatures of water volume while marginalizing the influence of less significant factors. We down-sample the synthetic data to a sampling frequency of 4 kHz, then generate five copies of the clean training and validation datasets, each corrupted with Gaussian noise according to the noise model in (5). The noise corrupted time domain data are transformed to the frequency domain according to model (7) and denoted as $\hat{X}_{\textrm{noised}}$ in the following. The real and imaginary components of $\hat{X}_{\textrm{noised}}$ are used as inputs to the network. We select $N_{f}=21$ frequencies between 30 to 90 Hz, see Fig. 4 as an example.
@@
The training process is implemented using the TensorFlow [2] and Keras [4] libraries. We employ the Adam optimizer [15] and utilize the RandomSearch algorithm from the Keras Tuner library [20] to optimize the model’s performance by exploring different hyperparameter configurations. The Keras Tuner is allowed to randomly select the activation function, learning rate, number of hidden layers, number of neurons per layer, and $L_{2}$ regularization penalty factor. The activation functions considered are “relu,” “sigmoid,” “tanh,” “selu,” “swish,” and “LeakyReLU,” while the number of hidden layers ranges from one to six. The number of neurons per layer varies between 100 and 5,000. The learning rate is selected from 1e-3, 1e-4, 1e-5, and the $L_{2}$ regularization penalty is chosen from 1e-5, 1e-6, 1e-7, 1e-8. For all networks in this study, the batch size is set to 256.
The training process is guided by minimizing the mean squared error (MSE) loss function, defined as
$\mathcal{L}(\theta)=\frac{1}{N_{{\rm train}}}\sum_{i=1}^{N_{{\rm train}}}\left(V_{{\rm true}}^{(i)}-\operatorname{NN}\left(\real{\hat{X}_{\textrm{noised}}^{(i)}},\mathbb{I}({\hat{X}_{\textrm{noised}}^{(i)}});\theta\right)\right)^{2}+\alpha\mathcal{R}(\theta),$ (8)
where $\operatorname{NN}$ denotes the neural network with weights and biases stored in $\theta$. Additionally, $N_{{\rm train}}$ represents the number of samples in the training dataset and $V_{{\rm true}}^{(i)}$ is the true water volume for the $i$-th sample. The term $\mathcal{R}(\theta)$ denotes the sum of the squared weights of the model and $\alpha$ is the $L_{2}$ regularization penalty coefficient. This loss function ensures that the predicted values closely approximate the true values. The model iteratively updates its weights and biases using the Adam optimizer to minimize this loss function. The convergence of the loss function over epochs is monitored to avoid overfitting. For a broader discussion on neural network training methodologies and optimization strategies, see [10, 24].
@@
The best-performing network is selected based on MSE for validation. After testing various hyperparameter combinations, the final network consists of five hidden layers with 2,570 (layer 1), 3,920 (layer 2), 3,360 (layer 3), 2,730 (layer 4), and 3,400 (layer 5) neurons. It employs the LeakyReLU activation function, a learning rate of 1e-5, and an $L_{2}$ regularization penalty factor of 1e-5. The output layer uses a linear activation function, and early stopping is enabled during both hyperparameter tuning and training to enhance generalization and prevent overfitting.
Figure 5 illustrates the approach for estimating water volume from seismic data. The figure shows the sequential steps in data processing, followed by the neural network model, which consists of five hidden layers. Additionally, it includes the learning curves for both training and validation, highlighting the model’s convergence and generalization during training.
Figure 5: Graphical illustration of the process for estimating water volume from seismic data. The left side illustrates the pre-processing phase, consisting deconvolution operation to normalize the source wavelet effects. On the right, the neural network-based estimation phase illustrates the applied network architecture. Each neuron in the hidden layers is represented by the symbol $a$, indicating its activation. In the architecture, $\{m_{1},m_{2},\ldots,m_{5}\}$ denote the number of neurons in each hidden layer. The loss curves as a function of epochs for the training and validation databases are shown in the top right corner.
@@ 5.1 Predictions of water volume
  • Field Measurements: Field data were acquired at seven distinct water table levels, ranging from a depth of –31.3 cm to –88.7 cm. At each level, three repeated measurements were recorded for each of three different drop heights, resulting in a dataset comprising $3\times 3\times 7=63$ observations. Using our knowledge of the actual water table level and the geometry of the sand pool, we can calculate the volume of the water-saturated zone. Additionally, based on a previous study [21], we assume the nominal porosity of the material to be 35 per cent.
@@
The Keras Tuner-optimized network architecture undergoes ten training runs, and the final water volume per sample result is determined as the average of these ten estimates. Figure 6 presents a comparison between the estimated (average) water volumes and their corresponding true values. The results for all drop heights are shown with different colors, red color denotes the repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). To get a crude approximation for the uncertainty, we assumed that the porosity value in the field measurements database is uncertain in a sense that we assumed $\phi_{\rm{true}}\in[0.95,1.05]\cdot 35$ per cent, that can be used to compute the error bars shown for each estimate with real data. The figure shows also the estimates for the synthetic dataset. These results demonstrate the potential of using proposed neural network-based approach to recover the water volume.
Figure 6: Left: Estimated volumes of water as a function of the true value. Red denotes repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green from drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue from drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). Right: Histogram of relative prediction errors, with the color coding of the asterisks corresponding to the same drop heights as shown on the left.
@@ 5.2 SHAP analysis
Next, we constructed two new receiver configurations and trained the neural network model for the field measurements based on SHAP values. For the first configuration, we selected ten receivers having the largest SHAP values, and for the second configuration, we randomly selected ten receivers from the full sensor array (see bottom panel of Fig. 7). The KerasTuner optimized network with SHAP analysis-based receiver selection consists three hidden layers with 3,490 (layer 1), 3,780 (layer 2), and 2,580 (layer 3) neurons, a LeakyRelu activation function, a learning rate of 1e-4, and $L_{2}$ regularization penalty factor of 1e-6. Similarly, the randomly selected receiver selection lead to optimized network with two hidden layers with 3,490 (layer 1) and 2,580 (layer 2) neurons, a LeakyRelu activation function, a learning rate of 1e-4, and $L_{2}$ regularization penalty factor of 1e-6.
@@
The normalized mean biases (NMB), mean absolute errors (MAE), and root mean square errors (RMSE) are used as evaluation metrics to quantitatively analyze the estimation results for all three receiver configurations, see Table 4. The table shows that the full receiver array results in the most accurate measures.
@@
Receiver array setup Database NMB (%) MAE (m3) RMSE (m3)